# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 2.2 Spatial error models with transformations
where $p_{\textbf{y}^{\ast}}(\cdot)$ is the density of $\textbf{y}^{\ast}$; see Section 2.1, and $\frac{dt_{\gamma}(y_{i})}{dy_{i}}$ is the derivative of the YJ transformation with respect to $y_{i}$; see Section S1 of the online supplement. The YJ-SEM with normally distributed errors (denoted as YJ-SEM-Gau) assumes that the error term in Equation (6) follows a normal distribution: $\mathbf{e}\sim N(0,\sigma_{\mathbf{e}}^{2}\mathbf{I}_{n})$. The YJ-SEM with Student’s t-distributed errors (denoted as YJ-SEM-t) assumes that the error term follows the structure specified in Equation (2). The log-likelihood function for y in YJ-SEM-Gau and YJ-SEM-t is:
@@
where for the YJ-SEM-Gau model, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the YJ-SEM-t model, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1. Note that $y_{i}=y^{\ast}_{i}$ for $i=1,...,n$, corresponds to an identity transformation in both SEM-Gau and SEM-t.
The three SEMs, SEM-Gau, SEM-t, and YJ-SEM-Gau, are special cases of the YJ-SEM-t model. The SEM-Gau model is a YJ-SEM-t model with the degrees of freedom parameter ($\nu$) set to $\infty$ and the YJ parameter ($\gamma$) set to 1, while the YJ-SEM-Gau model is a YJ-SEM-t model with $\nu=\infty$ and $\gamma\neq 1$. Finally, the SEM-t model is also a YJ-SEM-t model with $\gamma=1$.
Table 1: The definitions of r, M, and the parameters $\boldsymbol{\phi}$ for the four spatial error models. Each SEM has a mean vector of $\mathbf{X}\boldsymbol{\beta}$ and a covariance matrix of the form $\boldsymbol{\Sigma}=\sigma^{2}_{\textbf{e}}\textbf{M}^{-1}$, with $\textbf{A}=\textbf{I}_{n}-\rho\textbf{W}$.
Model $\boldsymbol{\phi}$ r M
SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu)^{\top}$ $\textbf{y}-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
YJ-SEM-Gau $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e},},\rho,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\textbf{A}$ A
YJ-SEM-t $(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$ $t_{\gamma}(\textbf{y})-\textbf{X}\boldsymbol{\beta}$ $^{\top}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{\tau}}\textbf{A}$ A
2.3 Spatial error models with missing responses
Consider that the response vector y of an SEM presented in Sections 2.1 and 2.2 contains missing values. Let $\textbf{y}_{o}$ be the subset of y with $n_{o}$ observed units, and $\textbf{y}_{u}$ be the subset of y with $n_{u}$ unobserved units. The complete response vector is $\textbf{y}=({\textbf{y}}_{o}^{\top},{\textbf{y}}_{u}^{\top})^{\top}$. A missing data indicator vector m of length ${n}$ containing 1’s and 0’s is defined. If an element in y is missing, then the corresponding element in m is 1 and 0, otherwise. In the presence of missing values, the vector r and the matrices X and M are decomposed into distinct components as follows:
@@ 2.3 Spatial error models with missing responses
where $\textbf{r}_{o}$ and $\textbf{r}_{u}$ are subvectors of r, corresponding to the observed and unobserved responses, respectively. Similarly, $\textbf{X}_{o}$ and $\textbf{X}_{u}$ represent the design matrices associated with the observed and unobserved responses. Additionally, the matrices $\textbf{M}_{oo}$, $\textbf{M}_{ou}$, $\textbf{M}_{uo}$, and $\textbf{M}_{uu}$ are submatrices of M.
We assume missing responses follow a missing not at random (MNAR) mechanism (Rubin,, 1976). The missing data mechanism is defined by the conditional distribution of m given y, commonly referred to as the missing data model, denoted by $p(\textbf{m}|\textbf{y},\boldsymbol{\psi},\textbf{X}^{*})$, where $\textbf{X}^{*}$ is an $n\times(q+1)$ design matrix containing the covariates of the missing data model. These covariates may be a subset of those used in SEMs. Assuming the conditional independence of m given y, $\textbf{X}^{*}$, and $\boldsymbol{\psi}$, the density is the product of $p(m_{i}\mid y_{i},\textbf{x}^{*}_{i},\boldsymbol{\psi})$ for $i=1,\ldots,n$, where $m_{i}$ and $y_{i}$ are the $i^{th}$ elements of m and y, and ${\textbf{x}_{i}^{*}}^{\top}$ is the $i^{th}$ row of $\textbf{X}^{*}$. The parameter vector $\boldsymbol{\psi}=(\boldsymbol{\psi}_{\textbf{x}}^{\top},\psi_{\textbf{y}})^{\top}$ consists of the fixed effects vector associated with covariates $\textbf{X}^{*}$, denoted as $\boldsymbol{\psi}_{\textbf{x}}=(\psi_{0},\psi_{1},\psi_{2},\ldots,\psi_{q})^{\top}$, and the fixed effect corresponding to y, denoted as $\psi_{\textbf{y}}$. A logistic regression model is used as the missing data model and is given by:
$p(\textbf{m}\mid\textbf{y},\textbf{X}^{*},\boldsymbol{\psi})=\prod_{i=1}^{n}\frac{e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\psi_{{y}}){m_{i}}}}{1+e^{{\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+\textbf{y}_{i}\psi_{\textbf{y}}}}.$ (10)
@@ 3 Variational Bayes inference for SEMs with missing data
Section 3.1 introduces the variational Bayes (VB) inference method. Section 3.2 discusses the efficient hybrid variational Bayes (HVB) methods for estimating the SEMs with missing data. Section 3.3 discusses the Deviance Information Criterion (DIC) to select the best SEMs. The VB method for the proposed SEMs in Section 2, when there are no missing data, is provided in Section S2 of the online supplement.
3.1 Variational Bayes inference
Variational Bayes (VB) inference approximates a Bayesian posterior distribution by formulating it as an optimisation problem, offering a more computationally efficient alternative to the computationally intensive Markov chain Monte Carlo (MCMC) methods for complex statistical models. This section presents the VB method for SEMs with missing data.
We assume that some values in the response vector of an SEM are missing under the missing not at random (MNAR) mechanism. Under this assumption, Bayesian estimation of the SEM requires consideration of the joint distribution of the missing data indicator vector m and the response vector y (Wijayawardhana et al.,, 2025), which is denoted by $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$. Using the selection model factorisation (Little and Rubin,, 2019), $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$ is expressed as
$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\textbf{y}\mid\boldsymbol{\xi}).$ (11)
where $p(\textbf{y}\mid\boldsymbol{\xi})$ is the likelihood of y. The log-likelihood for y for different SEMs is given in Section 2. Additionally, $p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$ denotes the density of the missing data model, given in Equation (10).
Consider Bayesian inference to estimate the joint posterior of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and the missing values $\textbf{y}_{u}$. For the SEM-Gau and YJ-SEM-Gau models, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t and YJ-SEM-t models, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$; see Table 1 for the definition of $\boldsymbol{\phi}$ for each SEM. Let the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ be denoted as $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$. We let $p(\boldsymbol{\xi})$ and $p(\boldsymbol{\psi})$ as the prior distributions of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, respectively. Using the selection model factorisation in Equation (11), the joint posterior distribution $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$ is given by
@@ 3.1 Variational Bayes inference
We define $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$. The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for each SEM with missing data, along with the total number of parameters to be estimated ($s_{m}$) for each model are provided in Table 2.
Table 2: The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$, and the total number of model parameters $s_{m}$ (including the length of the latent vector $\boldsymbol{\tau}$) for the SEM-Gau, YJ-SEM-Gau, SEM-t, and YJ-SEM-t models with missing values. The distribution $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$ represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2)).
Model $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ $s_{m}$
@@
$\mathcal{L}(\boldsymbol{\lambda})=E_{q}\left[\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\text{log}~{}q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})\right],$ (13)
where $E_{q}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$. Table 2 provides expressions for $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for different SEMs.
To maximise ELBO given in Equation (S6) with respect to variational parameters, $\boldsymbol{\lambda}$, stochastic gradient ascent (SGA) methods are used (Nott et al.,2012Titsias and Lázaro-Gredilla,20142015). The SGA method updates the initial value for $\boldsymbol{\lambda}$ (say $\boldsymbol{\lambda}^{(0)}$) according to the iterative scheme,
[…]

### v2

@@ 2.2 SEMs with transformations
[…]
The YJ-SEM with Gaussian errors (YJ-SEM-Gau) assumes that the error term in Equation (7) follows a normal distribution: $\textbf{e}\sim N(0,\sigma_{\textbf{e}}^{2}\textbf{I}_{n})$, where $\sigma_{\textbf{e}}^{2}$ is a variance parameter. The YJ-SEM with Student’s $t$-distributed errors (YJ-SEM-t) assumes that the error term follows the structure given in Equation (3). The log-likelihood function of y for YJ-SEM-Gau and YJ-SEM-t is expressed as:
@@
where for the YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the YJ-SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$. The definitions for r, M, and $\boldsymbol{\phi}$ are given in Table 1. Note that $y_{i}=y^{\ast}_{i}$ for i=1,,ni=1,...,n, corresponds to an identity transformation in both SEM-Gau and SEM-t.
The three SEMs, SEM-Gau, SEM-t, and YJ-SEM-Gau, are special cases of the YJ-SEM-t. The SEM-Gau is a YJ-SEM-t where the degrees of freedom parameter ($\nu$) equals $\infty$ and the YJ parameter ($\gamma$) equals 1, while the YJ-SEM-Gau is a YJ-SEM-t with $\nu=\infty$ and $\gamma\neq 1$. Finally, the SEM-t is also a YJ-SEM-t with $\gamma=1$.
3 SEMs with missing responses
This section develops the formulation of SEMs in settings where response values are missing and introduces the underlying missing data mechanisms, focusing in particular on the missing not at random (MNAR) mechanism.
Consider that the response vector y of an SEM presented in Sections 2.1 and 2.2 contains missing values. Let $\textbf{y}_{o}$ denote the subset of y with $n_{o}$ observed units, and $\textbf{y}_{u}$ denote the subset with $n_{u}$ unobserved (missing) units. The complete response vector can then be written as $\textbf{y}=(\textbf{y}_{o}^{\top},\textbf{y}_{u}^{\top})^{\top}$.
We define a binary missingness indicator vector $\textbf{m}=(m_{1},\ldots,m_{n})^{\top}$, where
$m_{i}=\begin{cases}1,&\text{if }y_{i}\text{ is missing,}\\ 0,&\text{if }y_{i}\text{ is observed.}\end{cases}$
In the presence of missing responses, the vector r and the matrices X and M, as defined in Table 1, are partitioned into corresponding observed and unobserved components as follows:
@@ 3 SEMs with missing responses
where $\textbf{r}_{o}$ and $\textbf{r}_{u}$ are subvectors of r corresponding to the observed and unobserved responses, respectively. Similarly, $\textbf{X}_{o}$ and $\textbf{X}_{u}$ represent the design matrices associated with the observed and unobserved responses, while $\textbf{M}_{oo}$, $\textbf{M}_{ou}$, $\textbf{M}_{uo}$, and $\textbf{M}_{uu}$ denote the corresponding submatrices of M.
3.1 Missing data mechanism and MNAR assumption
The missing data mechanism is characterised by the joint probability mass function of the missingness indicator vector $\textbf{m}=(m_{1},\ldots,m_{n})^{\top}$, where each element $m_{i}$ denotes whether the corresponding response $y_{i}$ is missing ($m_{i}=1$) or observed ($m_{i}=0$). Let $p(m_{i}\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})$ denote the probability that $y_{i}$ is missing, possibly depending on itself and on a set of covariates $\textbf{x}_{i}^{*}=(1,x_{i1},\ldots,x_{iq})^{\top}$. The parameter vector $\boldsymbol{\psi}=(\boldsymbol{\psi}_{\textbf{x}}^{\top},\psi_{\textbf{y}})^{\top}$ consists of the fixed effect parameter vector associated with covariates $\textbf{x}^{*}_{i}$, denoted as $\boldsymbol{\psi}_{\textbf{x}}=(\psi_{0},\psi_{1},\psi_{2},\ldots,\psi_{q})^{\top}$, and the fixed effect parameter corresponding to y, denoted as $\psi_{\textbf{y}}$.
Three main types of missing-data mechanisms are commonly distinguished following Rubin, (1976): (i) missing completely at random (MCAR), where the probability of missingness is independent of both the observed and unobserved data and on observed covariates, that is, $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})=p(m_{i}\mid\boldsymbol{\psi})$; (ii) missing at random (MAR), where the probability of missingness may depend on observed covariates but not on the observed and unobserved responses, so that $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})=p(m_{i}\mid\mathbf{x}_{i}^{*},\boldsymbol{\psi})$; and (iii) missing not at random (MNAR), where the probability of missingness depends on the response variable itself and observed covariates. Under MNAR, the term $p(m_{i}\mid y_{i},\mathbf{x}_{i}^{*},\boldsymbol{\psi})$ cannot be simplified further.
The conditional probability mass function of $m_{i}$ under MNAR, $p(m_{i}\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})$ can be modelled using any binary-response model, such as a logistic or probit regression. In this paper, we model the probability that the $i^{\text{th}}$ response is missing via a logistic regression:
$p(m_{i}=1\mid y_{i},\textbf{x}_{i}^{*},\boldsymbol{\psi})=\frac{\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})}{1+\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})},$ (11)
so that the joint probability mass function of vector m is
$p(\textbf{m}\mid\textbf{y},\textbf{X}^{*},\boldsymbol{\psi})=\prod_{i=1}^{n}\frac{\exp\{(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})m_{i}\}}{1+\exp(\textbf{x}_{i}^{*\top}\boldsymbol{\psi}_{\textbf{x}}+y_{i}\psi_{y})},$ (12)
where $\textbf{X}^{\ast}=(\textbf{x}^{\ast\top}_{1},\dots,\textbf{x}^{\ast\top}_{n})^{\top}$ denotes the design matrix for the missingness model, with dimension $n\times(q+1)$.
3.2 Joint modelling of y and m
In our study, the primary objective is the estimation of the parameters of the proposed SEMs, denoted by $\boldsymbol{\xi}$, which govern the distribution of the response vector y. However, under the MNAR mechanism, the missingness process is not ignorable, and valid inference requires modelling the joint distribution of $(\textbf{y},\textbf{m})$ or, equivalently, the joint estimation of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$ (Wijayawardhana et al.,, 2025). The joint distribution of y and m is denoted by $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$. Using the selection model factorisation (Little and Rubin,, 2019), $p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})$ can be factorised as
$p(\textbf{y},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})=p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\textbf{y}\mid\boldsymbol{\xi}),$ (13)
where $p(\textbf{y}\mid\boldsymbol{\xi})$ represents the likelihood of the SEM, and $p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})$ denotes the joint probability mass function of missingness indicators defined in Equation (12). The log-likelihood of $\mathbf{y}$ for the different SEMs, along with their full model specifications, is provided in Section 2.
It is important to note that the factorisation in Equation (13) is a standard and widely applicable decomposition for jointly modelling a missingness mechanism and a statistical model with missing data, and it applies to any SAR-type model; see Section S4 of the online supplement. This factorisation enables practical inference for SAR models with missing values under MNAR.  In Section 4, we discuss the estimation of the joint posterior distribution of the model parameters ($\boldsymbol{\xi},\boldsymbol{\psi}$) and $\textbf{y}_{u}$ using the proposed variational Bayes method.
@@ 4 Variational Bayes inference for SEMs with missing data
Section 4.1 provides an overview of the variational Bayes (VB) inference framework for estimating SEMs with missing responses. Section 4.2 then introduces an efficient hybrid variational Bayes (HVB) algorithm. For completeness, Section S2 of the online supplement presents the VB algorithm for SEMs with fully observed data.
4.1 Variational Bayes inference under missing responses
Variational Bayes (VB) inference approximates a posterior distribution by formulating it as an optimisation problem, offering a more computationally efficient alternative to the computationally intensive Markov chain Monte Carlo (MCMC) methods for complex statistical models. The standard VB approach for SEMs with missing responses is described below.
We assume that some values in the response vector of an SEM are missing under the MNAR mechanism, as described in Section 3. Under this assumption, Bayesian inference requires estimating the joint posterior distribution of the model parameters $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$, based on the joint likelihood defined in Equation (13). Since the subvector $\textbf{y}_{u}$ of $\textbf{y}=(\textbf{y}_{o}^{\top},\textbf{y}_{u}^{\top})^{\top}$ is unobserved, it must also be inferred. Consequently, Bayesian inference involves estimating the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$. For the SEM-Gau and YJ-SEM-Gau, $\boldsymbol{\xi}=\boldsymbol{\phi}$, while for the SEM-t and YJ-SEM-t, $\boldsymbol{\xi}=(\boldsymbol{\phi}^{\top},\boldsymbol{\tau}^{\top})^{\top}$; see Table 1 for the definition of $\boldsymbol{\phi}$ for each SEM.
Let the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ be denoted as $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$. We denote the prior distributions of $\boldsymbol{\xi}$ and $\boldsymbol{\psi}$ by $p(\boldsymbol{\xi})$ and $p(\boldsymbol{\psi})$, respectively. Using the selection model factorisation in Equation (13), the joint posterior distribution $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$} is given by
@@ 4.1 Variational Bayes inference under missing responses
We define $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$. The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for each SEM with missing responses, along with the total number of parameters to be estimated ($s$) for each model, are provided in Table 2.
Table 2: The term $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\xi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\xi})p(\boldsymbol{\psi})$, and the total number of model parameters $s$ (including the length of the latent vector $\boldsymbol{\tau}$ if exists) for the SEM-Gau, SEM-t, YJ-SEM-Gau, and YJ-SEM-t with missing responses. The density $p(\boldsymbol{\tau}\mid\boldsymbol{\phi})=\prod_{i=1}^{n}p(\tau_{i}\mid\nu)$} represents the $n$ independent inverse gamma latent variables used in SEMs with Student-$t$ errors (see Equation (2)).
Model $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ $s$
@@
We consider the variational approximation $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, indexed by the variational parameter $\boldsymbol{\lambda}$ to approximate the joint posterior $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$ in Equation (14). The VB approach approximates this posterior distribution by minimising the Kullback-Leibler (KL) divergence between $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ and $p(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m})$, which is defined as
@@
Minimising the KL divergence in Equation (S5) is equivalent to maximising the evidence lower bound (ELBO) on the logarithm of the marginal likelihood, $\text{log}~p(\textbf{y}_{o},\textbf{m})$, denoted by $\mathcal{L}(\boldsymbol{\lambda})$, where $p(\textbf{y}_{o},\textbf{m})=\int~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi})p\left(\boldsymbol{\xi},\boldsymbol{\psi}\right)d\boldsymbol{\boldsymbol{\xi}}d\boldsymbol{\psi}d\textbf{y}_{u}$ (Blei et al.,, 2017). The ELBO can be written as an expectation with respect to the variational distribution and is given by
$\mathcal{L}(\boldsymbol{\lambda})=\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\text{log}~q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})\right],$ (16)
where $\mathbb{E}_{{q}_{\boldsymbol{\lambda}}}\left[\cdot\right]$ denotes the expectation with respect to $q_{\boldsymbol{\lambda}}$. Table 2 provides expressions for $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ for different SEMs.
To maximise ELBO given in Equation (S6) with respect to variational parameters, $\boldsymbol{\lambda}$, the stochastic gradient ascent (SGA) method is used (Nott et al.,2012Titsias and Lázaro-Gredilla,20142015). The SGA method updates the initial value for $\boldsymbol{\lambda}$ (say $\boldsymbol{\lambda}^{(0)}$) according to the iterative scheme,
$\boldsymbol{\lambda}^{(t+1)}=\boldsymbol{\lambda}^{(t)}+\mathbcal{a}^{(t)}\circ\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda}^{(t)})}$, (17)
where $\widehat{\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})}$ is an unbiased estimate of the true gradient $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ of the ELBO in Equation (S6), $\mathbcal{a}^{(t)}$ ($t=0,1,\ldots$) is a vector of adaptive step sizes, obtained using the ADADELTA method (Zeiler,, 2012), described in Section S3 of the online supplement to facilitate rapid convergence of the SGA algorithm. The symbol $\circ$ represents the element-wise product of two vectors. The updating of Equation (17) is done until a stopping criterion is satisfied. Further details on the SGA method are provided in Section S2 of the online supplement.
In standard VB inference, the missing data vector, $\textbf{y}_{u}$, is treated as a set of parameters, and the joint posterior distribution of $\boldsymbol{\xi}$, $\boldsymbol{\psi}$, and $\textbf{y}_{u}$ is approximated using the joint variational distribution $q_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$. However, Wijayawardhana et al., (2025) demonstrated that this joint approximation is often insufficiently flexible to capture the posterior distributions of model parameters and $\textbf{y}_{u}$ simultaneously, due to the complex dependencies among the spatially correlated elements of $\textbf{y}_{u}$. To address this limitation, Section 4.2 introduces an alternative hybrid algorithm that effectively combines VB optimisation with MCMC sampling.